# Large Language Model with Region-guided Referring and Grounding for CT Report Generation

### v1

@@ Abstract
Computed tomography (CT) report generation is crucial to assist radiologists in interpreting CT volumes, which can be time-consuming and labor-intensive. Existing methods primarily only consider the global features of the entire volume, making it struggle to focus on specific regions and potentially missing abnormalities. To address this issue, we propose Reg2RG, the first region-guided referring and grounding framework for CT report generation, which enhances diagnostic performance by focusing on anatomical regions within the volume. Specifically, we utilize masks from a universal segmentation module to capture local features for each referring region. A local feature decoupling (LFD) strategy is proposed to preserve the local high-resolution details with little computational overhead. Then the local features are integrated with global features to capture inter-regional relationships within a cohesive context. Moreover, we propose a novel region-report alignment (RRA) training strategy. It leverages the recognition of referring regions to guide the generation of region-specific reports, enhancing the model’s referring and grounding capabilities while also improving the report’s interpretability. A large language model (LLM) is further employed as the language decoder to generate reports from integrated visual features, facilitating region-level comprehension. Extensive experiments on two large-scale chest CT-report datasets demonstrate the superiority of our method, which outperforms several state-of-the-art methods in terms of both natural language generation and clinical efficacy metrics while preserving promising interpretability. The code will be made publicly available.
@@ 1 Introduction
To reduce computational costs, previous works [5, 4] often downsample volumes or pool features. However, these methods may lose crucial high-resolution texture details [9]. To address this issue, we propose a local feature decoupling (LFD) strategy to preserve local high-resolution texture details and essential geometry with low computational cost. We further integrate global and local features to jointly capture significant inter-regional relationships within a cohesive context.
In medicine, interpretability is vital for radiologists to comprehend the visual basis of the generated reports. Previous works [10, 11] mainly use attention maps to link report keywords with image regions. However, these attention maps are often coarse and ambiguous, lacking precision and reliability. To enhance interpretability, we propose a novel region-report alignment (RRA) training strategy that enforces an explicit association between the referring region and the report. Specifically, the referring region order is shuffled for the model to identify anatomical areas before generating the report. By doing so, the reports are certainly grounded in the identified regions, enhancing interpretability and reliability.
@@ 3 Methods
In this section, we first overview our framework in Sec. 3.1. Next, we detail the local feature decoupling (LFD) strategy in Sec. 3.2, and explain their integration with global features in Sec. 3.3. Finally, we describe the region-report alignment (RRA) training strategy in Sec. 3.4.
@@ 3.1 Overview
The overview of our method Reg2RG is shown in Fig. 2. CT report generation involves creating report $\mathbf{R}$ based on CT volumes $\mathbf{V}\in\mathbb{R}^{H\times W\times D}$. Unlike the previous works [3, 5] that use only global features $\mathcal{G}$ of $\mathbf{V}$, we additionally extract a set of local features =$\mathcal{L}=\{\mathcal{L}_{1},\cdots,\mathcal{L}_{n}\}$ with the universal segmentation module. To preserve the local high-resolution texture details and significant geometry information with minimal computational overhead, local features are decoupled into texture and geometry. These local features work alongside global features to generate the report $\mathbf{R}$ using the LLM. Furthermore, we propose a training strategy that strengthens the correspondence between the local feature $\mathcal{L}_{i}$ and the region-specific report $R_{i}$, improving clarity and trustworthiness. The overall generating process can be described as follows:
@@ 3.2 Local Feature Decoupling Strategy
The region mask $M_{A_{j}}$ is then utilized to construct the corresponding local features $\mathcal{L}_{A_{j}}$. To preserve higher-resolution details without increasing much computational burden, we design a local feature decoupling (LFD) strategy that separates texture and geometry information.
For the texture information, we first use region mask $M_{A_{j}}$ to extract the region volume $V_{A_{j}}$ from $\mathbf{V}$ by element-wise multiplication. This results in a large redundant area outside the target region, which is not informative. Therefore, we crop the region volume $V_{A_{j}}$ to exclude these irrelevant parts. Since $V_{A_{j}}$ is much smaller than the entire volume $\mathbf{V}$, we can retain higher-resolution details without increasing the input size. Next, we utilize a 3D volume encoder $f_{V}$ to extract local texture features $\mathcal{L}_{A_{j}}^{t}$. The following adapter module $f_{A}$ is used to compress and align these local texture features with the embedding space of the LLM. This process can be formulated as follows:
@@
The previous works [25, 26, 27, 23, 24] solely consider texture features, ignoring the essential geometry features for assessing lesion size and location in medical contexts. In contrast, our method includes it as a supplementary feature. Specifically, we introduce the geometry information by encoding the region mask $M_{A_{j}}$, which is uncropped to preserve the original size and position. A lightweight volume encoder $f_{M}$ is used to extract geometry features $\mathcal{L}_{A_{j}}^{g}$, followed by a fully connected layer $f_{P}$ to project these features for LLM input. This process is formulated as follows:
@@ 3.3 Global-Local Features Collaboration
In medicine, different regions are interactive rather than isolated. Therefore, we include global features for context. Specifically, the same volume encoder $f_{V}$ and adapter $f_{A}$ (as in Sec. 3.2) are used to extract global features $\mathcal{G}$:
@@ 3.4 Region-Report Alignment Training Strategy
While the segmentation module can provide the anatomical area name, this information may bias the model during report generation since it can lead the model to rely on the name rather than the region features. Therefore, we train the model to recognize this information independently from the region features. This strategy helps the model better understand the referred region, making the generated report more reliably grounded. Furthermore, we shuffle the order of the local features at each step to prevent the model from identifying anatomical areas by a fixed sequence. Thus, although the content of $\mathcal{L}$ and $\mathcal{L}_{A}$ is unchanged, the correspondence of each region feature varies. The generation process is defined as follows:
@@
During training, region-level reports $R_{i}$ of each referring region are used as ground truth. In addition, we add a prefix $P_{i}$=“The region [$i$is [area name]” to indicate the anatomical area name. The restructured report remains a sequence of text tokens: $\mathbf{R}=\{(P_{1},R_{1}),\cdots,(P_{n},R_{n})\}=\{r_{1},r_{2},\cdots,r_{T}\}$ ($T$ is the report length). This format integrates seamlessly with the native auto-regressive training process of the LLM. The training process is optimized by minimizing the language-modeling loss as shown below:
@@
During inference, the prefix $P_{i}$ is removed from the generated report $\mathbf{R}$ before evaluation.
@@ 4.2 Implementation Details and Baselines
We use the text-prompted segmentation model SAT [6] to extract anatomical masks when unavailable. Both global and local volumes are resized to 256 $\times$ 256 $\times$ 64 for input. We use the pre-trained ViT3D and Perceiver from RadFM [4] as the volume encoder $f_{V}$ and adapter $f_{A}$ to extract texture features, each represented by 32 visual embeddings. Given that geometric information is sparser than texture information, a lightweight 3-layer ViT3D serves as the mask encoder $f_{M}$ to extract geometric features. These features are pooled into a single embedding and projected by $f_{P}$ to align with the LLM’s embedding space. For the LLM, we choose LLaMA2-7B [38] and utilize LoRA [39] for parameter-efficient fine-tuning. We train the model with the AdamW optimizer [40] at an initial learning rate of $5\times 10^{-5}$, following a constant learning rate schedule with a warmup phase. Training on the RadGenome-Chest dataset takes 48 hours on two RTX 4090 GPUs with PyTorch 2.0, running for 6 epochs with an effective batch size of 16. For the CTRG-Chest-584K dataset, training takes 24 hours for 10 epochs with the same batch size. To optimize memory usage, ZeRO [41] stage 2 is applied alongside gradient checkpointing [42]. We use the checkpoint of the last epoch for evaluation.
@@ 4.4 Ablation Study
To demonstrate the effectiveness of each component in our model, we conduct comprehensive ablation studies on the RadGenome-ChestCT dataset [16]. Given that our volume encoder $f_{V}$ and adapter $f_{A}$ are initialized using the pre-trained checkpoint from RadFM [4], we take it as the baseline for comparison. First, we examine the contribution of the local feature decoupling (LFD) strategy in Sec. 4.4.1. Next, we investigate the effectiveness of global-local features collaboration by incrementally adding different visual features in Sec. 4.4.2. Additionally, we analyze the performance of our region-report alignment (RRA) training strategy in Sec. 4.4.3. Finally, we validate the necessity of employing large-scale LLMs for region-level referring and grounding in Sec. 4.4.4. The results of these experiments are detailed in Table 4, Table 5 and Table 6, demonstrating the individual and collective impact of each component.
@@ 4.4.1 Local Feature Decoupling Strategy
Table 5: The effectiveness of different visual features and region-report alignment (RRA) training strategy. The best results are highlighted in bold. The TXT and GEO denote the texture and geometric information of local features, respectively, whereas GLB stands for the global features.
Setting Visual Features RRA CE Metrics NLG Metrcis
@@ 4.4.3 Region-Report Alignment Training Strategy
In setting (d) of Table 5, we validate the efficacy of our region-report alignment (RRA) training strategy, which guides the LLM to generate reports based on referring region information for reliable grounding. The results indicate that our complete model outperforms the setting (c) model across all CE and NLG metrics except ROUGE-L. This improvement demonstrates the effectiveness of the RRA strategy in aligning region-specific features with report generation, ensuring more accurate and clinically relevant outputs. The slightly lower ROUGE-L score can be attributed to its reliance on exact sequence matching, which may overlook the use of synonyms or paraphrased expressions. As a result, ROUGE-L may not fully reflect report quality in cases where the phrasing differs but the semantic content remains consistent. The top METEOR performance supports this, as it better captures semantic similarity by considering synonyms and paraphrasing.
@@ 4.5.2 Analysis of Generated Reports
4.5.3 Report with Referring and Grounding
A key advantage of our model is its ability to generate reports explicitly grounded to specific regions. As illustrated in Fig. 5, our generated reports are segmented into distinct sections, each corresponding to a particular anatomical region. We use the same color scheme as in Fig. 4 to represent the region masks and reports associated with different regions.
Each regional report begins with the referring region, providing an initial hint about the focus of the LLM. The area recognition result is presented before the report, helping verify the reliability of the generated report. Successfully identifying the referring region indicates that the LLM refers to the correct regional information, enabling the report to be properly grounded in the corresponding region. This enhances the model’s interpretability and reliability of the reports, which is valuable for clinical practice. Conversely, if the referenced region is misidentified, the report becomes unreliable regardless of its content and cannot be definitely linked to any region. This strategy provides a straightforward yet effective mechanism for validating reports and offers reliable reference information to assist radiologists in their interpretation.
Figure 4: Case studies of our model and the SOTA MedVInT [22]. The gray background highlights incorrect diagnoses.
Figure 5: Report grounding cases of our model. The different colors correspond to distinct anatomical regions.

### v2

@@ Abstract
Computed tomography (CT) report generation is crucial to assist radiologists in interpreting CT volumes, which can be time-consuming and labor-intensive. Existing methods primarily only consider the global features of the entire volume, making it struggle to focus on specific regions and potentially missing abnormalities. To address this issue, we propose Reg2RG, the first region-guided referring and grounding framework for CT report generation, which enhances diagnostic performance by focusing on anatomical regions within the volume. Specifically, we utilize masks from a universal segmentation module to capture local features for each referring region. A local feature decoupling (LFD) strategy is proposed to preserve the local high-resolution details with little computational overhead. Then the local features are integrated with global features to capture inter-regional relationships within a cohesive context. Moreover, we propose a novel region-report alignment (RRA) training strategy. It leverages the recognition of referring regions to guide the generation of region-specific reports, enhancing the model’s referring and grounding capabilities while also improving the report’s interpretability. A large language model (LLM) is further employed as the language decoder to generate reports from integrated visual features, facilitating region-level comprehension. Extensive experiments on two large-scale chest CT-report datasets demonstrate the superiority of our method, which outperforms several state-of-the-art methods in terms of both natural language generation and clinical efficacy metrics while preserving promising interpretability. The code is available at https://github.com/zhi-xuan-chen/Reg2RG.
@@ 1 Introduction
To reduce computational costs, previous works [5, 4] often downsample volumes or pool features. However, these methods may lose crucial high-resolution texture details [9]. To address this issue, we propose a local feature decoupling (LFD) strategy to preserve local high-resolution texture details and essential geometry with low computational cost.
Since the global volume-level embedding provides significant inter-regional relationships within a cohesive context, we further integrate global features with local features to enable collaboration between them. However, achieving this effective global-local collaboration is non-trivial. Since global features encompass information from all regions, they may interfere with the diagnosis of a specific region, as validated in Table 6. To address this challenge, we propose a novel region-report alignment (RRA) training strategy that enhances the alignment between local features and their corresponding region reports, reducing the influence of irrelevant global information and ensuring that the generated reports are more robust and accurate.
In medicine, interpretability is vital for radiologists to comprehend the visual basis of the generated reports. Previous works [10, 11] mainly use attention maps to link report keywords with image regions. However, these attention maps are often coarse and ambiguous, lacking precision and reliability. The proposed RRA training strategy can also enhance interpretability by establishing a clear association between the referring region and the report. This explicit alignment ensures that the generated reports are firmly grounded in the identified regions, thereby improving both interpretability and reliability.
@@ 3 Methods
In this section, we first overview our framework in Sec. 3.1. Next, we detail the local feature decoupling (LFD) strategy in Sec. 3.2, and explain its integration with global features in Sec. 3.3. Finally, we describe the region-report alignment (RRA) training strategy in Sec. 3.4.
@@ 3.1 Overview
The overview of our method Reg2RG is shown in Fig. 2. CT report generation involves creating report $\mathbf{R}$ based on CT volumes $\mathbf{V}\in\mathbb{R}^{H\times W\times D}$. Unlike the previous works [3, 5] that use only global features $\mathcal{G}$ of $\mathbf{V}$, we additionally extract a set of local features =$\mathcal{L}=\{\mathcal{L}_{1},\cdots,\mathcal{L}_{n}\}$ with the universal segmentation module. To preserve the local high-resolution texture details and significant geometry information with minimal computational overhead, local features are decoupled into texture and geometry. The decoupled local features work alongside global features to generate the report $\mathbf{R}$ using the LLM. Furthermore, we propose a training strategy that strengthens the alignment between the local feature $\mathcal{L}_{i}$ and the region-specific report $R_{i}$, thereby improving diagnostic accuracy and enhancing the reliability of the generated report. The overall generating process can be described as follows:
@@ 3.2 Local Feature Decoupling Strategy
The region mask $M_{A_{j}}$ is then utilized to construct the corresponding local features $\mathcal{L}_{A_{j}}$. To preserve higher-resolution details without increasing much computational burden, we design a local feature decoupling (LFD) strategy that separates texture and geometry information. Texture information refers to patterns within regions of interest that capture surface characteristics, which are essential for disease diagnosis. For example, in lung CT scans, the texture feature “ground-glass opacity” indicates increased lung tissue density. Geometry information encompasses the size, shape, and spatial location of regions of interest, playing a crucial role in medical imaging. For example, the geometry feature “enlarged heart” may indicate conditions such as heart failure or pericardial effusion.
For the texture information, we first use region mask $M_{A_{j}}$ to extract the region volume $V_{A_{j}}$ from $\mathbf{V}$ by element-wise multiplication. This results in a large redundant area outside the region of interest, which is not informative. Therefore, we crop the region volume $V_{A_{j}}$ to exclude these irrelevant parts. Since $V_{A_{j}}$ is much smaller than the entire volume $\mathbf{V}$, we can retain higher-resolution details without increasing the input size. Next, we utilize a 3D volume encoder $f_{V}$ to extract local texture features $\mathcal{L}_{A_{j}}^{t}$. The following adapter module $f_{A}$ is used to compress and align these local texture features with the embedding space of the LLM. This process can be formulated as follows:
@@
Previous works [25, 26, 27, 23, 24] extract local features by only encoding cropped regions, focusing exclusively on texture features while neglecting the essential geometry features necessary for assessing lesion size and location in medical contexts. In contrast, our method incorporates them as supplementary features. Specifically, we introduce the geometry information by encoding the region mask $M_{A_{j}}$, which is uncropped to preserve the original size and position. A lightweight mask encoder $f_{M}$ is used to extract geometry features $\mathcal{L}_{A_{j}}^{g}$, followed by a fully connected layer $f_{P}$ to project these features for LLM input. This process is formulated as follows:
@@ 3.3 Global-Local Features Collaboration
In medicine, different regions are interrelated rather than isolated. Therefore, we incorporate global features to provide contextual information. Specifically, the same volume encoder $f_{V}$ and adapter $f_{A}$ (as in Sec. 3.2) are used to extract global features $\mathcal{G}$:
@@ 3.4 Region-Report Alignment Training Strategy
While the segmentation module can provide the anatomical area name, this information may introduce bias during report generation, as the model might rely on the name rather than the actual local features. For instance, if the input includes the name “lung” for a region, the model may focus solely on generating content related to the “lung”, potentially overlooking the real information within the local features. Therefore, we train the model to recognize this information independently from the local features. This strategy helps the model better understand the referred region, making the generated report more reliably grounded. Furthermore, we shuffle the order of the local features $\mathcal{L}_{A}$ at each step to prevent the model from associating anatomical areas with a fixed sequence. Thus, although the content of $\mathcal{L}$ and $\mathcal{L}_{A}$ remains unchanged, the order of each local feature varies between them. The generation process is defined as follows:
@@
During training, region-level reports $R_{i}$ of each referring region are used as ground truth. In addition, we add a prefix $P_{i}$=“The region [$i$is [area name]” before $R_{i}$ to indicate the anatomical area name. The model learns to recognize the anatomical area name by predicting this prefix. Since anatomical region names are fixed for each specific region, they provide a clear and structured target for the model, enabling it to develop a more precise semantic understanding of the referred regions. The restructured report remains a sequence of text tokens: $\mathbf{R}=\{(P_{1},R_{1}),\cdots,(P_{n},R_{n})\}=\{r_{1},r_{2},\cdots,r_{T}\}$ ($T$ is the report length). This format integrates seamlessly with the native auto-regressive training process of the LLM. The training process is optimized by minimizing the language-modeling loss as shown below:
@@
When evaluating the quality of the generated report, the prefix $P_{i}$ is removed from the $\mathbf{R}$.
@@ 4.2 Implementation Details and Baselines
We use the text-prompted segmentation model SAT-Pro [6] to extract anatomical masks when unavailable. The SAT-Pro model is configured with 256 object queries, and its input CT intensities are normalized to the range [0, 1] using min-max normalization. For our input, both global and local volumes are resized to 256 $\times$ 256 $\times$ 64. We use the pre-trained ViT3D and Perceiver from RadFM [4] as the volume encoder $f_{V}$ and adapter $f_{A}$ to extract texture features, each represented by 32 visual embeddings. Given that geometric information is sparser than texture information, a lightweight 3-layer ViT3D serves as the mask encoder $f_{M}$ to extract geometric features. These features are pooled into a single embedding and projected by $f_{P}$ to align with the LLM’s embedding space. For the LLM, we choose LLaMA2-7B [38] and utilize LoRA [39] for parameter-efficient fine-tuning. The LoRA configuration is set with a rank $r=8$, a scaling factor $\alpha=32$, and a dropout rate of 0.1.
We train the model with the AdamW optimizer [40] at an initial learning rate of $5\times 10^{-5}$, following a constant learning rate schedule with a warmup phase. Training on the RadGenome-Chest dataset takes 48 hours on two RTX 4090 GPUs with PyTorch 2.0, running for 6 epochs with an effective batch size of 16. For the CTRG-Chest-584K dataset, training takes 24 hours for 10 epochs with the same batch size. To optimize memory usage, ZeRO [41] stage 2 is applied alongside gradient checkpointing [42]. We use the checkpoint of the last epoch for evaluation.
@@ 4.3.3 Region Recognition Results
4.3.4 Region-level Reports Evaluation
Table 4 showcases the evaluation results of the region-level reports generated by our model on the RadGenome-ChestCT dataset [16]. We observe that the performance varies for reports from different anatomical regions. The higher occurrence frequency of certain regions in reports contributes to better performance on the abdomen, mediastinum, heart, and lung. In contrast, performance tends to be lower for regions that appear less frequently, such as the breast and thyroid. The subpar CE performance on the bone and trachea & bronchi may be attributed to the RadBERT model from CT-CLIP [15] rarely considering abnormalities in these regions. Investigating approaches to improve the diagnosis of less common regions and abnormalities could serve as a direction for future research.
Table 4: The evaluation of region-level reports generated by our model. The Tra. & Bro. region refers to the trachea and bronchi. The Amount represents the number of reports that include descriptions for the specific region.
Region Amount CE Metrics NLG Metrcis
Precision Recall F1-score BLEU-1 BLEU-4 METEOR ROUGE-L
Abdomen 1517 0.546 0.288 0.377 39.50 25.13 43.33 38.52
Bone 1509 0.000 0.000 0.000 41.05 28.49 47.15 42.13
Breast 56 0.174 0.211 0.190 22.46 13.63 23.14 21.96
Esophagus 1323 0.222 0.021 0.039 59.58 45.16 56.39 57.49
Heart 1418 0.320 0.135 0.190 39.85 27.44 44.75 43.19
Lung 1514 0.393 0.126 0.191 27.22 15.51 31.74 30.82
Mediastinum 1513 0.557 0.242 0.337 37.65 20.09 41.98 33.68
Pleura 1169 0.333 0.069 0.115 37.20 26.91 38.84 47.21
Thyroid 42 0.125 0.040 0.061 21.83 10.64 19.49 19.47
Tra. & Bro. 1401 0.000 0.000 0.000 48.42 38.93 61.47 56.18
@@ 4.4 Ablation Study
To demonstrate the effectiveness of each component in our model, we conduct comprehensive ablation studies on the RadGenome-ChestCT dataset [16]. Given that our volume encoder $f_{V}$ and adapter $f_{A}$ are initialized using the pre-trained checkpoint from RadFM [4], we take it as the baseline for comparison. First, we examine the contribution of the local feature decoupling (LFD) strategy in Sec. 4.4.1. Next, we investigate the effectiveness of global-local features collaboration by incrementally adding different visual features in Sec. 4.4.2. Additionally, we analyze the performance of our region-report alignment (RRA) training strategy in Sec. 4.4.3. Finally, we validate the necessity of employing large-scale LLMs for region-level referring and grounding in Sec. 4.4.4. The results of these experiments are detailed in Table 5, Table 6, Table 7 and Table 8, demonstrating the individual and collective impact of each component.
@@ 4.4.1 Local Feature Decoupling Strategy
Table 6: The effectiveness of different visual features. The best results are highlighted in bold. The TXT and GEO denote the texture and geometric information of local features, respectively, whereas GLB stands for the global features.
Setting Visual Features CE Metrics NLG Metrcis
@@ 4.4.3 Region-Report Alignment Training Strategy
In Table 7, we validate the efficacy of our region-report alignment (RRA) training strategy, which guides the LLM to generate reports based on referring region information for reliable grounding. The results indicate that our model with RRA outperforms the one without it across all CE and NLG metrics, except for ROUGE-L. This improvement demonstrates the effectiveness of the RRA strategy in aligning region-specific features with report generation, ensuring more accurate and clinically relevant outputs. The slightly lower ROUGE-L score can be attributed to its reliance on exact sequence matching, which may overlook the use of synonyms or paraphrased expressions. As a result, ROUGE-L may not fully reflect report quality in cases where the phrasing differs but the semantic content remains consistent. The top METEOR performance supports this, as it better captures semantic similarity by considering synonyms and paraphrasing.
@@ 4.4.4 LLM as the Language Decoder
Table 7: The effectiveness of region-report alignment (RRA) strategy.
RRA CE Metrics NLG Metrcis
Precision Recall F1-score BLEU-1 BLEU-4 METEOR ROUGE-L
0.372 0.176 0.239 46.21 23.53 42.94 37.01
0.423 0.181 0.253 47.25 24.87 44.07 36.65
@@ 4.5.2 Analysis of Generated Reports
4.5.3 Region-level Reports with Referring and Grounding
A key advantage of our model is its ability to generate region-level reports explicitly grounded to specific regions. As illustrated in Fig. 5, our generated reports are segmented into distinct sections, each corresponding to a particular anatomical region. We use the same color scheme as in Fig. 4 to represent the region masks and reports associated with different regions.
Each region-level report begins with the referring region, providing an initial hint about the focus of the LLM. The area recognition result is presented before the report, helping verify the reliability of the generated report. Successfully identifying the referring region indicates that the LLM refers to the correct regional information, enabling the report to be properly grounded in the corresponding region. This enhances the model’s interpretability and reliability of the reports, which is valuable for clinical practice. Conversely, if the referenced region is misidentified, the report becomes unreliable regardless of its content and cannot be definitely linked to any region. This strategy provides a straightforward yet effective mechanism for validating reports and offers reliable reference information to assist radiologists in their interpretation.
Figure 4: Case studies of our model and the SOTA MedVInT [22]. The different colors represent distinct anatomical areas, as shown at the bottom of each example. The gray background highlights incorrect diagnoses.
Figure 5: Region-level reports generated by our model. Each regional report refers to a specific region and is grounded in the anatomical area depicted in the left figure. The different colors correspond to distinct anatomical regions.
4.5.4 Segmentation Results on the Two Datasets
We present two segmentation cases from both the RadGenome-ChestCT and CTRG-Chest-584K datasets in Fig. 6. Benefiting from the superior performance of SAT [6], the segmentation results on the RadGenome-ChestCT dataset are highly satisfactory, providing accurate regional information that enables our model to achieve outstanding performance. Regarding the CTRG-Chest-584K dataset, its segmentation results are relatively coarser due to the lower quality of the volumes in this dataset. This may limit the potential of our model, as evidenced by Table 1, where its improvement over the second-best method on CTRG-Chest-584K is smaller than that on RadGenome-ChestCT. However, our model still outperforms other methods by effectively leveraging regional information, demonstrating the robustness and effectiveness of our framework.
Figure 6: Segmentation results on the RadGenome-ChestCT and CTRG-Chest-584K datasets. The different colors represent distinct anatomical areas, as shown at the bottom.
5 Limitations and Future Directions
Despite the encouraging outcomes, several limitations of our model warrant further investigation in the future.
Our framework utilizes the segmentation results to acquire local features. The inaccurate or incomplete masks, caused by segmentation failures, could potentially affect the performance of our model to some extent. Therefore, strengthening the robustness of our method against segmentation errors is valuable for improving the reliability of the generated reports. Given that the relative positions of anatomical regions are generally consistent across patients, incorporating anatomical spatial priors can help refine the anatomical masks. This prior anatomical knowledge enhances the model’s robustness to segmentation failures, thereby ensuring the quality of the extracted local features. Moreover, we plan to explore integrating segmentation model optimization into the training process to further enhance the quality of the anatomical masks, thereby better supporting diagnosis.
Another limitation is the insufficiency of detailed lesion information. The current local features in our framework are restricted to organ-level information, which is inadequate for precise lesion detection. As shown in Fig. 4, the model may miss certain lesions or misidentify their locations. To enhance the model’s ability to detect and characterize lesions, we plan to explore integrating lesion segmentation or detection into the framework. Incorporating lesion-specific information will enable the model to capture more detailed insights into abnormalities, thereby facilitating the generation of more accurate and clinically relevant reports.